OpenStudy (anonymous):

During a hike, Vijay saw a total of 19 birds and butterflies. Knowing that birds have 2 wings and butterflies have 4 wings, he counted 52 wings. Use a model to determine how many birds he saw.

OpenStudy (anonymous):

choices: 9 10 11 12

OpenStudy (diyadiya):

let the number of birds be x and Butterflies be y x + y = 19 ( since he saw a total of 19 of them) Birds have 2 wings & butterflies have 4 wings we can say that 2 times Number of Birds plus 4 times number of Butterflies is equal to 52 2x + 4y = 52 Now we have two equations x+y=19 2x+4y = 52 can you solve these two equations?

OpenStudy (anonymous):

Let the birds be x and butterflies be y. then , x+y=19 , y=19-x and , 2x+4y=52 2x+4(19-x)=52 2x+76-4x=50 -2x=50-76 -2x=-26 x=-26/-2 x=13, birds=13